OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

Please Help...

OpenStudy (phi):

parallel means the same slope. what is the slope of y = x + 7 ? to find the slope, you match your equation with y = m x + b m is the slope

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

y+2=1(x-6) y+2=x-6 y=x-4

OpenStudy (anonymous):

oh so you added the pair into the equation